write an equation in the form y = mx + b for the line which contains the points (-8,20) and (-4,0).
Answer by rothauserc(4718)

In y = mx +b, m is the slope of the line
m = (0 - 20) / (-4 - (-8)) = -20 / 4 = -5
now we have y = -5x +b
we calculate b by using one of the points, pick (-8, 20)
20 = -5(-8) +b
20 = 40 +b
b = -20
therefore the equation of this line is
y = -5x -20